PHY3

Please if you are not sure 100% dont touch it in order not get bad rate. You will get Thump Up for correct answer.4. +-14 points TR4 3.P.017 Calculate Amax for blackbody radiation for the following (a) liquid helium (2.5 K) (b) room temperature (293 K) (c) a steel furnace (1450 K) (d) a blue star (8000 K)
Answer #1

4] Use Wein's displacement law relating maximum wavelength emitted and the temperature.

XmarT_ 2.898 × 10-3 mK

a]

So, for T = 2.5 K [liquid helium]

\lambda_{max} = \frac{2.898\times 10^{-3} }{2.5} = 1159.2 \mu m

b] T = 293 K,

\lambda_{max} = \frac{2.898\times 10^{-3} }{293} = 9.89 \mu m

c] T = 1450 K,

\lambda_{max} = \frac{2.898\times 10^{-3} }{1450} = 1.998 \mu m

d] T = 8000 K,

\lambda_{max} = \frac{2.898\times 10^{-3} }{8000} = 0.362 \mu m.